OpenStudy (aravindg):

In a triangle ABC,if [1/(a+c)]+[1/(b+c)]=3/(a+b+c) then c is equal to??

OpenStudy (aravindg):

a,b, c are angles

OpenStudy (aravindg):

?can u help ash

OpenStudy (aravindg):

@satellite73 , @AccessDenied , @campbell,

OpenStudy (anonymous):

A, B, C are the angles and a, b, c are the sides? this is standard....

OpenStudy (aravindg):

ys i made a a mistake of using small letters thats why i told they are angles

OpenStudy (anonymous):

i guess a start is \[\frac{1}{a+c}+\frac{1}{b+c}=\frac{1}{120}\]

OpenStudy (aravindg):

180 !!

OpenStudy (anonymous):

oops right

OpenStudy (anonymous):

\[\frac{1}{a+c}+\frac{1}{b+c}=\frac{1}{60}\]

OpenStudy (anonymous):

we can easily get rid of one of the variables, but i don't know how to get rid of two of them

OpenStudy (kinggeorge):

Does c have to be a specific number? Or can it be given in terms of a and b?

OpenStudy (aravindg):

my options are a)30 b)60 c)75 d)90

OpenStudy (kinggeorge):

By method of mostly guess and check, answer b is correct. A 60-60-60 triangle satisfies the relationship.

OpenStudy (aravindg):

any other logical explanation?

jhonyy9 (jhonyy9):

- so i think the first step will be 1/(a+c) +1/(b+c) = 3/(a+b+c) = 3/180 = 1/60 - so than 1/(a+c) +1/(b+c) = 1/60 - if a+b+c=180 so than a=180-b-c - than 1/(180-b-c+c) +1/(b+c) = 1/60 1/(180-b) +1/(b+c) =1/60 - how you think please this is one right way or ... ?

OpenStudy (kinggeorge):

You can simplify your original equation to\[c = {(-a b+60 a-b^2+120 b)\over(a+b-60)}\]From there, I used wolfram alpha to let \(c=30, 60, 75, 90\) until I got a solution in integers that fit the relationship.

OpenStudy (aravindg):

THANKS A LOT SEE IF U CAN HELP HERE http://openstudy.com/study?login#/updates/4f81e8dee4b0505bf0837aff

OpenStudy (kinggeorge):

For that matter however, a=40, b=65, c=75 also works.

OpenStudy (kinggeorge):

So answer c. could also be correct :/